OpenStudy (anonymous):

If f(x)=2x+6 and g(x)=x^3, what is (g。f)(0)?

OpenStudy (anonymous):

The dot in the middle means we are looking for g( f(0) ). So, lets figure out f(0). Since f(x)=2x+6 then f(0)=0+6=6. Now we find g(6). Since g(x)=x^3 then g(6)=6^3=216

OpenStudy (anonymous):

Okie dokie

OpenStudy (anonymous):

216 is the answer?

OpenStudy (anonymous):

yep